### v1

# Foundation Model-Aided Hierarchical Control for Robust RIS-Assisted Near-Field Communications

Abstract
The deployment of extremely large aperture arrays (ELAAs) in sixth-generation (6G) networks could shift communication into the near-field communication (NFC) regime. In this regime, signals exhibit spherical wave propagation, unlike the planar waves in conventional far-field systems. Reconfigurable intelligent surfaces (RISs) can dynamically adjust phase shifts to support NFC beamfocusing, concentrating signal energy at specific spatial coordinates. However, effective RIS utilization depends on both rapid channel state information (CSI) estimation and proactive blockage mitigation, which occur on inherently different timescales. CSI varies at millisecond intervals due to small-scale fading, while blockage events evolve over seconds, posing challenges for conventional single-level control algorithms. To address this issue, we propose a dual-transformer (DT) hierarchical framework that integrates two specialized transformer models within a hierarchical deep reinforcement learning (HDRL) architecture, referred to as the DT-HDRL framework. A fast-timescale transformer processes ray-tracing data for rapid CSI estimation, while a vision transformer (ViT) analyzes visual data to predict impending blockages. In HDRL, the high-level controller selects line-of-sight (LoS) or RIS-assisted non-line-of-sight (NLoS) transmission paths and sets goals, while the low-level controller optimizes base station (BS) beamfocusing and RIS phase shifts using instantaneous CSI. This dual-timescale coordination maximizes spectral efficiency (SE) while ensuring robust performance under dynamic conditions. Simulation results demonstrate that our approach improves SE by approximately 18% compared to single-timescale baselines, while the proposed blockage predictor achieves an F1-score of 0.92, providing a 769 ms advance warning window in dynamic scenarios.
I INTRODUCTION
The evolution toward sixth-generation (6G) wireless systems will be driven by the demand for unprecedented data rates, ultra-low latency, and massive connectivity [saad2019vision]. A key enabler is the deployment of extremely large aperture arrays (ELAAs) operating in millimeter-wave (mmWave) and terahertz (THz) bands, which offer the massive spatial multiplexing gain required to meet 6G targets [rappaport2019wireless]. In high-frequency bands, small wavelengths enable practical deployment of ELAAs with hundreds of antenna elements. However, the combination of large apertures and high frequencies requires a paradigm shift from the far-field planar wave assumptions to the near-field communication (NFC) regime [dai2022near]. In NFC, signals propagate as spherical waves, making channel characteristics dependent on both angle and distance to the receiver [an2023codebook]. Given that the NFC in 6G systems can extend beyond 100 meters [bjornson2022massive], strictly angular beamsteering is insufficient. Instead, beamfocusing is required to precisely concentrate signal energy at specific spatial locations.
Realizing the potential of NFC is impeded by two primary physical constraints: severe propagation loss and spherical wavefront complexity. The mmWave and THz bands are susceptible to severe path loss and are easily blocked by physical objects [giordani2019survey]. Moreover, spherical wavefronts introduce significant nonlinearity into channel state information (CSI) estimation, requiring models that account for spatial variations across the large array aperture [cui2023near, he2023near]. This makes conventional CSI acquisition computationally prohibitive due to high pilot overhead [heath2016overview], yet accurate high-dimensional CSI is a prerequisite for effective beamfocusing [ma2021sparse].
Reconfigurable Intelligent Surfaces (RISs) have emerged as a transformative solution to mitigate propagation losses and mitigate blockages [direnzo2020smart]. Comprising massive arrays of passive elements, RISs manipulate incident signal phases to establish virtual LoS links, effectively reconfiguring the electromagnetic environment [wu2020towards, zhang2022beam]. However, RIS deployment increases the CSI estimation challenge by introducing cascaded channel links, while simultaneously requiring proactive control to respond to blockage events. This leads to a dual-timescale control problem: CSI must be estimated at millisecond intervals to track fast fading, while blockage prediction operates on second-scale dynamics as objects move through the environment.
I-A Related Work
I-A1 NFC Channel Modeling and Estimation
The transition to ELAAs has necessitated a re-evaluation of channel models. Unlike far-field systems that assume planar wavefronts, near-field channels depend on both angle and distance, significantly increasing estimation complexity [tang2023nearfield]. Recent works have explored phenomena such as beam squint in wideband systems [zhang2023beamsquint]. However, accurately estimating these high-dimensional channels remains a bottleneck. While deep learning approaches have been suggested [wang2024transformer, elbir2020deep], traditional CNN-based estimators struggle with the long-range spatial dependencies inherent to large aperture arrays. Furthermore, few existing works explicitly leverage the geometric priors of near-field propagation to reduce computational complexity in transformer architectures, a gap our work addresses.
I-A2 Blockage Prediction
Beyond channel estimation, link blockage poses a critical challenge. Traditional approaches such as relay switching are reactive. Recent research has shifted toward proactive strategies using multi-modal sensors [alkhateeb2018machine, charan2021vision]. Methods fusing multi-modal data have shown promise in building environment-aware networks [ghassemi2024multi]. Vision transformers (ViTs) demonstrate exceptional performance in blockage prediction by processing sequential image data [wang2022vision, ghassemi2025generative]. ViTs identify motion patterns of users and objects that indicate future blockages, offering superior generalization compared to classical signal processing methods [sanchez2022millimeter]. By integrating predictive capabilities, a network can preemptively reconfigure resources to maintain connectivity.
I-A3 Hierarchical Control in Wireless Networks
Reinforcement learning (RL) has been widely applied to RIS optimization [ghassemi2025foundation]. However, RL agents suffer from the problem of dimensionality when controlling ELAAs. Hierarchical deep reinforcement learning (HDRL) offers a solution by decomposing timescales [huang2023hierarchical, liu2021deep]. Such decomposition is critical for RIS-assisted systems where beamforming must adapt to fast fading while phase shifts may be updated on a slower frame-based schedule [deng2022reconfigurable]. Our work extends this by decoupling the problem based on the information dynamics of CSI and blockage prediction results.
I-B Motivations and Contributions
To address these challenges, we propose a dual-transformer (DT) hierarchical framework that integrates specialized transformer models within a HDRL architecture for proactive control in RIS-assisted NFC systems. Our approach coordinates proactive blockage prediction with real-time CSI estimation through a dual-timescale design that handles the inherent mismatch between slow blockage dynamics and fast channel fading. Recent advances in RIS-assisted systems have demonstrated the potential of intelligent reflecting surfaces for enhancing wireless coverage and capacity [liu2024ris], motivating our integration of foundation models with hierarchical control.
The primary contributions are summarized as follows:
  • Multi-modal Transformer Design: We design two specialized transformer models tailored for the NFC regime. The first processes ray-tracing data to extract spatial features for CSI estimation. The second acts as a predictive module, utilizing a ViT to analyze sequential visual data and predict future blockage dynamics.
  • Dual-Timescale HDRL Control: We introduce a hierarchical control mechanism where a high-level agent leverages slow-timescale blockage predictions for strategic mode selection (LoS vs. RIS-assisted non-line-of-sight (NLoS)), while a low-level agent utilizes fast-timescale CSI for real-time joint beamfocusing and RIS optimization problem to maximize the sum spectral efficiency (SE) based on these estimates.
  • Robustness and Efficiency: We validate the proposed framework using the high-fidelity DeepVerse 6G O1 ray-tracing dataset. The evaluation includes a comprehensive sensitivity analysis across varying ELAA dimensions and RIS element counts, alongside a rigorous assessment of computational complexity to verify the real-time feasibility of the dual-timescale design under dynamic blockage conditions.
II System Model and Problem Formulation
In this section, we develop the mathematical model for RIS-assisted NFC and formulate the joint beamfocusing problem for the base station (BS) and RIS. We explicitly account for the spherical wave propagation inherent to ELAAs and establish the optimization variables that will be handled by the AI agents.
II-A System Model
We consider a downlink NFC multiple-input single-output (MISO) system comprising a BS equipped with an ELAA, a passive RIS, and $K$ single-antenna user equipment (UEs) indexed by $\mathcal{K}=\{1,...,K\}$. Each UE is located at $\mathbf{u}_{k}=[x_{k},y_{k},z_{k}]^{T}$ in Cartesian coordinates.
II-A1 Base Station Configuration
The BS is equipped with an ELAA configured as a uniform planar array (UPA) of $N=N_{y}\times N_{z}$ antennas aligned along the $y$- and $z$-axes. The array center is located at the origin $\mathbf{u}_{\text{BS}}=[0,0,0]^{T}$. Adopting linear indexing $n\in\{1,\dots,N\}$, the position of the $n$-th antenna element relative to the array center is given by:
$\mathbf{p}_{n}^{\text{BS}}=[0,n_{y}d_{\text{BS}},n_{z}d_{\text{BS}}]^{T},$ (1)
where $d_{\text{BS}}$ is the inter-element spacing (in meters), $n_{y}\in\mathcal{N}_{y}$ and $n_{z}\in\mathcal{N}_{z}$ are the local indices along the $y$- and $z$-axes, respectively, and $\cdot]^{T}$ denotes the transpose operator. To ensure the array is centered at $\mathbf{u}_{\text{BS}}$, the local indices take values from the sets $\mathcal{N}_{y}=\{-\frac{N_{y}-1}{2},\dots,\frac{N_{y}-1}{2}\}$ and $\mathcal{N}_{z}=\{-\frac{N_{z}-1}{2},\dots,\frac{N_{z}-1}{2}\}$, respectively.
II-A2 Reconfigurable Intelligent Surface Configuration
The RIS is a UPA composed of $M=M_{y}\times M_{z}$ passive metasurface elements. The RIS center is located at $\mathbf{u}_{\text{RIS}}=[x_{\text{RIS}},y_{\text{RIS}},z_{\text{RIS}}]^{T}$. Similar to the BS, the position of the $m$-th metasurface element relative to the RIS center is defined as:
$\mathbf{p}_{m}^{\text{RIS}}=[0,m_{y}d_{\text{RIS}},m_{z}d_{\text{RIS}}]^{T},$ (2)
where $d_{\text{RIS}}$ is the RIS element spacing, and the indices $m_{y},m_{z}$ are drawn from the centered sets $\mathcal{M}_{y}=\{-\frac{M_{y}-1}{2},\dots,\frac{M_{y}-1}{2}\}$ and $\mathcal{M}_{z}=\{-\frac{M_{z}-1}{2},\dots,\frac{M_{z}-1}{2}\}$. The absolute global coordinate of the $m$-th element is $\mathbf{u}_{\text{RIS}}+\mathbf{p}_{m}^{\text{RIS}}$. The metasurface response is characterized by the reflection coefficient matrix $\bm{\Theta}=\text{diag}(e^{j\phi_{1}},...,e^{j\phi_{M}})\in\mathbb{C}^{M\times M}$, where $\phi_{m}\in[0,2\pi)$ is the controllable phase shift of the $m$-th element.
II-A3 NFC Spherical Wave Channel Model
We adopt the spherical wave model which accounts for both phase variations and distance-dependent path loss across the array aperture [bjornson2022massive].
1) BS-to-UE Direct Channel: The distance from the $n$-th BS antenna to the $k$-th UE is $d_{k,n}^{\text{BD}}=\|\mathbf{u}_{k}-(\mathbf{u}_{\text{BS}}+\mathbf{p}_{n}^{\text{BS}})\|=\|\mathbf{u}_{k}-\mathbf{p}_{n}^{\text{BS}}\|$, where $\|\cdot\|$ denotes the Euclidean norm. The channel coefficient from the $n$-th BS antenna to UE $k$ is modeled as [bjornson2022massive]:
$h_{\text{BD},k}]_{n}=\sqrt{\eta_{k,n}^{\text{BD}}}\cdot e^{-j\frac{2\pi}{\lambda}d_{k,n}^{\text{BD}}},$ (3)
where $\lambda$ is the carrier wavelength and $\eta_{k,n}^{\text{BD}}=(\frac{\lambda}{4\pi d_{k,n}^{\text{BD}}})^{2}$ is the free-space path loss. This model assumes a deterministic line-of-sight channel with spherical wave propagation. The full channel vector is denoted by $\mathbf{h}_{\text{BD},k}\in\mathbb{C}^{N\times 1}$.
2) RIS-to-UE Channel: Similarly, the distance from the $m$-th metasurface element to the $k$-th UE is $d_{k,m}^{\text{RD}}=\|\mathbf{u}_{k}-(\mathbf{u}_{\text{RIS}}+\mathbf{p}_{m}^{\text{RIS}})\|$. The $m$-th element of the channel vector $\mathbf{h}_{\text{RD},k}\in\mathbb{C}^{M\times 1}$ is given by:
$h_{\text{RD},k}]_{m}=\sqrt{\eta_{k,m}^{\text{RD}}}\cdot e^{-j\frac{2\pi}{\lambda}d_{k,m}^{\text{RD}}},$ (4)
where $\eta_{k,m}^{\text{RD}}=(\frac{\lambda}{4\pi d_{k,m}^{\text{RD}}})^{2}$ represents the path loss for the RIS-to-UE link, and $\cdot]_{m}$ denotes the $m$-th element of a vector.
3) BS-to-RIS Channel: The channel between the BS and RIS is denoted by $\mathbf{G}_{\text{BR}}\in\mathbb{C}^{M\times N}$, where the $(m,n)$-th entry is:
$\mathbf{G}_{\text{BR}}]_{m,n}=\sqrt{\eta_{m,n}^{\text{BR}}}\cdot e^{-j\frac{2\pi}{\lambda}d_{m,n}^{\text{BR}}},$ (5)
here, $d_{m,n}^{\text{BR}}=\|(\mathbf{u}_{\text{RIS}}+\mathbf{p}_{m}^{\text{RIS}})-(\mathbf{u}_{\text{BS}}+\mathbf{p}_{n}^{\text{BS}})\|$ represents the distance between the $n$-th BS antenna and $m$-th metasurface element, $\eta_{m,n}^{\text{BR}}=(\frac{\lambda}{4\pi d_{m,n}^{\text{BR}}})^{2}$ is the corresponding path loss, and $\cdot]_{m,n}$ denotes the $(m,n)$-th element of a matrix.
II-B NFC Propagation Analysis
We analyze the phase discrepancy between near-field and far-field assumptions. The boundary between these regions is defined by the Rayleigh distance $Z_{R}$ [bjornson2022massive, dai2022near]:
$Z_{R}=\frac{2D^{2}}{\lambda},$ (6)
where $D$ is the maximum array aperture dimension. For ELAAs, $D$ is large relative to $\lambda$, causing $Z_{R}$ to extend significantly. For instance, with $N=1024$ antennas configured as $32\times 32$ at $d_{\text{BS}}=\lambda/2$ and $f_{c}=3.5$ GHz, the aperture $D\approx 1.88$ m yields $Z_{R}\approx 82$ m, placing typical urban deployment distances firmly in the near-field regime. This analysis applies generally to any ELAA (BS or RIS). For a user located at $\mathbf{u}_{k}$ with radial distance $r=\|\mathbf{u}_{k}\|$ and angle $\theta$ relative to the array boresight, let $q_{n}$ denote the generic scalar position of the $n$-th element along the array aperture. The phase of the signal under the planar wave assumption is $\phi_{\text{planar}}(n)=\frac{2\pi}{\lambda}q_{n}\sin(\theta)$. However, the exact spherical phase is determined by the Euclidean distance [bjornson2022massive]:
$\phi_{\text{spherical}}(n)=\frac{2\pi}{\lambda}\left(\sqrt{r^{2}+q_{n}^{2}-2rq_{n}\sin(\theta)}-r\right).$ (7)
Using a second-order Taylor expansion, the phase difference (error) is approximated as:
$\Delta\phi(n)\approx\frac{\pi q_{n}^{2}\cos^{2}(\theta)}{\lambda r}.$ (8)
A related challenge in wideband near-field systems is beam squint, where the beam direction varies across frequency subcarriers due to the wavelength-dependent phase term $e^{-j\frac{2\pi}{\lambda}d}$ [gao2023wideband]. Since $\lambda=c/f$ varies with frequency, each subcarrier in an OFDM system experiences a different phase shift, causing the focused beam to deviate spatially across the signal bandwidth. In the near-field regime, where distances $d_{k,n}$ vary significantly across the array aperture, this frequency-dependent beam deviation becomes severe, leading to signal distortion and power loss. Traditional far-field beamforming cannot compensate for this effect, as it assumes angle-only dependence without accounting for the distance-frequency coupling inherent to spherical wavefronts.
For distances $r<Z_{R}$, this phase error $\Delta\phi(n)$ becomes significant (exceeding $\pi/8$), causing destructive interference if conventional far-field beamsteering is used. As we demonstrate in Section III, our CSI transformer architecture explicitly accounts for this non-linear phase term via exact Cartesian distance feature inputs $d_{k,n}=\|\mathbf{u}_{k}-\mathbf{p}_{n}^{\text{BS}}\|$, enabling precise energy focusing and implicit beam squint compensation through learned distance-frequency relationships.
II-C Problem Formulation
We define a binary blockage indicator $b_{k}\in\{0,1\}$ for UE $k$, where $b_{k}=1$ denotes an available LoS link. To maximize efficiency while minimizing complexity, we assume a switching strategy where each UE is served either via the direct link or the RIS-assisted link.
Let $\mathbf{W}=[\mathbf{w}_{1},...,\mathbf{w}_{K}]\in\mathbb{C}^{N\times K}$ be the BS beamfocusing matrix and $\mathbf{s}=[s_{1},...,s_{K}]^{T}\in\mathbb{C}^{K\times 1}$ be the symbol vector with $\mathbb{E}[|s_{k}|^{2}]=1$ for all $k\in\mathcal{K}$. The unified received signal at UE $k$ is:
$y_{k}=\mathbf{h}_{\text{eff},k}^{H}\mathbf{w}_{k}s_{k}+\sum_{i\neq k}\mathbf{h}_{\text{eff},k}^{H}\mathbf{w}_{i}s_{i}+n_{k},$ (9)
where $n_{k}\sim\mathcal{CN}(0,\sigma^{2})$ is the additive white Gaussian noise (AWGN) at UE $k$ with variance $\sigma^{2}$, and the effective channel vector is defined as:
$\mathbf{h}_{\text{eff},k}^{H}=b_{k}\mathbf{h}_{\text{BD},k}^{H}+(1-b_{k})\mathbf{h}_{\text{RD},k}^{H}\bm{\Theta}\mathbf{G}_{\text{BR}},$ (10)
where $\mathbf{h}_{\text{BD},k}\in\mathbb{C}^{N\times 1}$ is the direct BS-to-UE channel vector for user $k$ with elements defined in Eq. (3), $\mathbf{h}_{\text{RD},k}\in\mathbb{C}^{M\times 1}$ is the RIS-to-UE channel vector for user $k$ with elements defined in Eq. (4), $\mathbf{G}_{\text{BR}}\in\mathbb{C}^{M\times N}$ is the BS-to-RIS channel matrix with elements defined in Eq. (5), and $(\cdot)^{H}$ denotes the Hermitian operator.
The signal-to-interference-plus-noise ratio (SINR) is given by:
$\text{SINR}_{k}=\frac{|\mathbf{h}_{\text{eff},k}^{H}\mathbf{w}_{k}|^{2}}{\sum_{i\neq k}|\mathbf{h}_{\text{eff},k}^{H}\mathbf{w}_{i}|^{2}+\sigma^{2}}.$ (11)
The spectral efficiency (SE) for UE $k$ is defined as:
$\text{SE}_{k}=\log_{2}(1+\text{SINR}_{k})\quad\text{(bps/Hz)},$ (12)
where $\text{SINR}_{k}$ is given in Eq. (11).
Optimization with Estimated Inputs: In practical scenarios, the network controller cannot access the perfect channel states or future blockage events instantaneously. Instead, it must rely on the predicted blockage state $\hat{b}_{k}$ (provided by the ViT) and the estimated channel $\hat{\mathbf{h}}_{\text{eff},k}$ (provided by the CSI transformer).
The predicted SE for UE $k$ is given by:
$\widehat{\text{SE}}_{k}=\log_{2}(1+\widehat{\text{SINR}}_{k}),$ (13)
where $\widehat{\text{SINR}}_{k}$ is computed using the estimated channel $\hat{\mathbf{h}}_{\text{eff},k}$ and predicted blockage state $\hat{b}_{k}$. Consequently, we formulate the joint beamfocusing and reflection optimization problem to maximize the sum SE based on these estimates:
$\max_{\mathbf{W},\bm{\Theta}}$ $\sum_{k=1}^{K}\widehat{\text{SE}}_{k}$ (14)
[…]
$|e^{j\phi_{m}}|=1,\quad\forall m\in\{1,.,M\},\quad\text{(C2)}$
$\widehat{\text{SE}}_{k}\geq\text{SE}_{\text{min}},\quad\forall k\in\mathcal{K},\quad\text{(C3)}$
where $\widehat{\text{SE}}_{k}$ is defined in Eq. (13), $||\cdot||_{F}$ denotes the Frobenius norm, $P_{\text{max}}$ is the maximum transmit power budget at the BS, and $\text{SE}_{\text{min}}$ is the minimum required spectral efficiency per user to ensure quality of service (QoS). Constraint (C1) enforces the transmit power limitation at the BS. Constraint (C2) ensures that each RIS element maintains unit-modulus reflection, which is a physical requirement of passive metasurfaces. Constraint (C3) guarantees that each user achieves the minimum QoS requirement.
Problem (14) is non-convex due to the multiplicative coupling of $\mathbf{W}$ and $\bm{\Theta}$ in the effective channel and the unit-modulus constraints in (C2). Solving this via exhaustive search implies exponential complexity with respect to $M$, while alternating optimization methods are too slow for real-time constraints. Furthermore, the inherent timescale disparity—where blockages evolve slower than fast fading—makes it inefficient to update both CSI and blockage predictions at the same rate. This necessitates a hierarchical decomposition into long-term prediction and short-term adaptation.
III Proposed Dual-Transformer Hierarchical Framework
To solve the non-convex optimization problem in (14), we propose a hierarchical framework, referred to as the Dual-Transformer HDRL (DT-HDRL) framework, that leverages two specialized transformer architectures with an HDRL agent, as illustrated in Fig. 1.
Figure 1: System architecture showing dual transformer inputs (CSI transformer from ray-tracing simulations and ViT from camera sensors) feeding into HDRL with high-level strategic planning and low-level beamfocusing optimization.
III-A Transformer-Based NFC CSI Estimation
We deploy a transformer-based foundation model designed to extract latent representations of the NFC channel structure, capturing the long-range spatial dependencies inherent to large-aperture arrays.
III-A1 Input Feature Extraction
For each UE $k$ and BS antenna element $n$, we extract three geometric features from the environment that characterize the propagation path. We assume that the BS has access to UE location information $\mathbf{u}_{k}=[x_{k},y_{k},z_{k}]^{T}$, which can be obtained through wireless data. Based on the known UE locations, we compute:
  • Distance: $d_{k,n}=\|\mathbf{u}_{k}-\mathbf{p}_{n}^{\text{BS}}\|$ (meters)
  • Elevation angle: $\vartheta_{k,n}=\arcsin\left(\frac{z_{k}-z_{n}^{\text{BS}}}{d_{k,n}}\right)$ (radians)
  • Azimuth angle: $\psi_{k,n}=\arctan\left(\frac{y_{k}-y_{n}^{\text{BS}}}{x_{k}-x_{n}^{\text{BS}}}\right)$ (radians).
The input sequence for UE $k$ is then constructed as:
$\mathbf{x}_{k}=\{[d_{k,n},\vartheta_{k,n},\psi_{k,n}]\}_{n=1}^{N}\in\mathbb{R}^{N\times 3}.$ (15)
III-A2 Feature Embedding and Positional Encoding
The geometric inputs $\mathbf{x}_{k}\in\mathbb{R}^{N\times 3}$ are projected into a latent space of dimension $d_{\text{model}}$. To preserve the spatial topology of the UPA, we add learnable positional encodings $\mathbf{P}\in\mathbb{R}^{N\times d_{\text{model}}}$. The input to the first encoder layer is $\mathbf{H}^{(0)}=\text{Linear}(\mathbf{x}_{k})+\mathbf{P}$, where $\text{Linear}(\cdot):\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}^{N\times d_{\text{model}}}$ denotes a learnable linear projection layer implemented as $\text{Linear}(\mathbf{x}_{k})=\mathbf{x}_{k}\mathbf{W}_{\text{proj}}+\mathbf{b}_{\text{proj}}$, with weight matrix $\mathbf{W}_{\text{proj}}\in\mathbb{R}^{3\times d_{\text{model}}}$ and bias vector $\mathbf{b}_{\text{proj}}\in\mathbb{R}^{d_{\text{model}}}$.
III-A3 Multi-Head Self-Attention (MHSA) Formulation
The MHSA mechanism [vaswani2017attention] enables the model to capture non-local spatial correlations across the ELAA by allowing each antenna element to attend to all other elements simultaneously. Given the input $\mathbf{H}^{(l)}\in\mathbb{R}^{N\times d_{\text{model}}}$ from layer $l$, for each attention head $i\in\{1,\dots,h\}$, we compute Query $\mathbf{Q}_{i}\in\mathbb{R}^{N\times d_{k}}$, Key $\mathbf{K}_{i}\in\mathbb{R}^{N\times d_{k}}$, and Value $\mathbf{V}_{i}\in\mathbb{R}^{N\times d_{v}}$ matrices via linear projections of the input $\mathbf{H}^{(l)}$:
$\mathbf{Q}_{i}$ $=\mathbf{H}^{(l)}\mathbf{W}_{i}^{Q},\quad\mathbf{W}_{i}^{Q}\in\mathbb{R}^{d_{\text{model}}\times d_{k}},$ (16)
$\mathbf{K}_{i}$ $=\mathbf{H}^{(l)}\mathbf{W}_{i}^{K},\quad\mathbf{W}_{i}^{K}\in\mathbb{R}^{d_{\text{model}}\times d_{k}},$ (17)
$\mathbf{V}_{i}$ $=\mathbf{H}^{(l)}\mathbf{W}_{i}^{V},\quad\mathbf{W}_{i}^{V}\in\mathbb{R}^{d_{\text{model}}\times d_{v}},$ (18)
where $\mathbf{W}_{i}^{Q},\mathbf{W}_{i}^{K},\mathbf{W}_{i}^{V}$ are learnable weight matrices that project the input representation into different subspaces. The attention output for head $i$ is computed as:
$\text{Head}_{i}=\text{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}_{i}\in\mathbb{R}^{N\times d_{v}},$ (19)
where all antenna elements attend to each other without restrictions, enabling the model to capture the full spatial correlation structure of the near-field channel. Following the standard transformer architecture [vaswani2017attention], the outputs from all $h$ heads are concatenated and projected through a learnable output weight matrix $\mathbf{W}^{O}\in\mathbb{R}^{hd_{v}\times d_{\text{model}}}$:
$\text{MHSA}(\mathbf{H}^{(l)})=\text{Concat}(\text{Head}_{1},\dots,\text{Head}_{h})\mathbf{W}^{O}\in\mathbb{R}^{N\times d_{\text{model}}}$
This mechanism dynamically weighs the contribution of each antenna element based on the spherical wavefront curvature, prioritizing elements with constructive phase coherence.
III-A4 Output Projection and Training
The final layer output is aggregated via mean pooling and passed through a linear projection head to produce the estimated effective channel:
$\hat{\mathbf{h}}_{\text{eff},k}=\mathbf{W}_{\text{out}}\left(\frac{1}{N}\sum_{n=1}^{N}\mathbf{h}_{k,n}^{(L)}\right)+\mathbf{b}_{\text{out}}\in\mathbb{C}^{N\times 1},$ (21)
where $L$ is the number of transformer encoder layers, $\mathbf{h}_{k,n}^{(L)}\in\mathbb{R}^{d_{\text{model}}}$ denotes the output of the $L$-th layer for the $n$-th antenna element and user $k$, $\mathbf{W}_{\text{out}}\in\mathbb{C}^{N\times d_{\text{model}}}$ is the learnable output weight matrix, and $\mathbf{b}_{\text{out}}\in\mathbb{C}^{N\times 1}$ is the bias vector that maps the aggregated features to complex-valued channel coefficients. The model is trained to minimize the mean squared error (MSE) between predicted and ground-truth channels:
[…]
where $\mathcal{L}_{\text{MSE}}$ denotes the MSE loss function used for training the CSI transformer. Under the assumption of Gaussian channel statistics, this MSE objective provides an efficient approximation to maximum likelihood estimation [kay1993fundamentals], enabling the transformer to learn optimal channel representations from limited training data.
III-B ViT-Based Blockage Prediction
We employ a ViT to serve as a predictive module for environmental dynamics, leveraging its ability to capture the temporal motion dependencies that precede blockage events.
III-B1 Patch Embedding and Linear Projection
The input consists of $F$ frames, each of resolution $H\times W$ with $C$ channels. We flatten each frame into a sequence of 2D patches $\mathbf{x}_{p}\in\mathbb{R}^{N_{p}\times(P^{2}\cdot C)}$, where $P$ is the patch size and $N_{p}=HW/P^{2}$ is the number of patches. To handle the temporal dimension, we stack frames channel-wise, resulting in an effective channel dimension $C^{\prime}=F\times C$. Each patch is then represented as a flattened vector of dimension $P^{2}\cdot C^{\prime}$, and the complete input sequence has dimension $N_{p}\times(P^{2}\cdot C^{\prime})$, where $N_{p}=HW/P^{2}$ is the number of patches. The patches are linearly projected into a latent vector space of dimension $D$:
$\mathbf{z}_{0}=[\mathbf{x}_{\text{class}};\mathbf{x}^{1}_{p}\mathbf{E};\mathbf{x}^{2}_{p}\mathbf{E};\dots;\mathbf{x}^{N_{p}}_{p}\mathbf{E}]+\mathbf{E}_{pos},$ (23)
where $\mathbf{E}\in\mathbb{R}^{(P^{2}C^{\prime})\times D}$ is the learnable patch embedding matrix, and $\mathbf{E}_{pos}\in\mathbb{R}^{(N_{p}+1)\times D}$ represents the learnable positional embeddings required to retain spatial information. The prepended $\mathbf{x}_{\text{class}}$ is the learnable classification token whose state at the output of the transformer serves as the blockage prediction representation.
III-B2 Blockage Probability Output
The sequence $\mathbf{z}_{0}\in\mathbb{R}^{(N_{p}+1)\times D}$ passes through $L$ layers of MHSA and Multi-Layer Perceptron (MLP) blocks, where each row represents a token (either a patch embedding or the class token). The final prediction is obtained by passing the class token state $\mathbf{z}^{L}_{0}$ through a sigmoid-activated head [dosovitskiy2020image]:
$\hat{\mathbf{p}}=\delta(\text{MLP}(\text{LayerNorm}(\mathbf{z}^{L}_{0}))),$ (24)
where $\hat{\mathbf{p}}=[\hat{p}_{1},\dots,\hat{p}_{K}]^{T}\in[0,1]^{K}$ is the vector of blockage probabilities. Here, $\hat{p}_{k}$ represents the likelihood that the LoS link for user $k$ will be blocked in the next macro-interval. This probability captures the environmental uncertainty and serves as a key state input for the HDRL meta-controller. The ViT is trained via supervised learning to minimize the binary cross-entropy (BCE) loss [dosovitskiy2020image]:
$\mathcal{L}_{\text{BCE}}=-\frac{1}{K}\sum_{k=1}^{K}\Big[b_{k}\log(\hat{p}_{k}+\epsilon)+(1-b_{k})\log(1-\hat{p}_{k}+\epsilon)\Big],$ (25)
where $b_{k}\in\{0,1\}$ is the ground-truth blockage label for user $k$ derived from ray-tracing data as described in Section IV-A, $\mathcal{L}_{\text{BCE}}$ denotes the BCE loss function, and $\epsilon=10^{-7}$ is a small constant added for numerical stability to prevent logarithm of zero.
1Input: Ray-tracing dataset $\mathcal{D}_{ray}$, Visual dataset $\mathcal{D}_{vis}$, Number of episodes $E$, Macro-steps per episode $T_{\text{macro}}$, Micro-steps per macro-step $N_{\text{macro}}$, Batch size $B$, Learning rate $\eta$, Discount factors $\gamma_{l}$ (sub-controller) and $\gamma_{h}$ (meta-controller), Target network update rate $\tau$;
2 Output: Trained CSI transformer $\theta_{csi}$, ViT $\theta_{vit}$, HDRL policies $\pi_{h},\pi_{l}$;
3 Initialize: CSI transformer $\theta_{csi}$ with random weights;
4 Initialize: ViT $\theta_{vit}$ with random weights;
5 Initialize: Meta-agent policy $\pi_{h}$ and Q-network $Q_{h}$ with random weights and their target networks $\pi_{h}^{\prime}$, $Q_{h}^{\prime}$;
6 Initialize: Sub-agent policy $\pi_{l}$ and Q-network $Q_{l}$ with random weights and their target networks $\pi_{l}^{\prime}$, $Q_{l}^{\prime}$;
7 Initialize: Replay buffers $\mathcal{B}_{h}\leftarrow\emptyset$ (capacity $|\mathcal{B}_{h}|$), $\mathcal{B}_{l}\leftarrow\emptyset$ (capacity $|\mathcal{B}_{l}|$);
8 Initialize: Ornstein-Uhlenbeck noise process with standard deviation $\sigma$ for exploration;
9
// Phase 1: Foundation Model Pre-training. Train CSI transformer and ViT independently using supervised learning
// Train CSI Transformer for channel estimation
10 Sample batch $\{(\mathbf{x}^{(i)},\mathbf{h}_{\text{eff}}^{(i)})\}_{i=1}^{B}$ from $\mathcal{D}_{ray}$;
11 Forward pass: $\hat{\mathbf{h}}_{\text{eff}}^{(i)}\leftarrow\text{Transformer}_{\theta_{csi}}(\mathbf{x}^{(i)})$ for $i=1,\ldots,B$;
[…]
13 Backward pass and update: $\theta_{csi}\leftarrow\theta_{csi}-\eta\nabla_{\theta_{csi}}\mathcal{L}_{\text{MSE}}$;
14
// Train ViT for blockage prediction
15 Sample batch $\{(img^{(i)},\mathbf{b}_{label}^{(i)})\}_{i=1}^{B}$ from $\mathcal{D}_{vis}$;
16 Forward pass: $\hat{\mathbf{p}}^{(i)}\leftarrow\text{ViT}_{\theta_{vit}}(img^{(i)})$ for $i=1,\ldots,B$;
17 Compute BCE loss $\mathcal{L}_{\text{BCE}}$ via Eq. (25) with numerical stability constant $\epsilon$;
18 Backward pass and update: $\theta_{vit}\leftarrow\theta_{vit}-\eta\nabla_{\theta_{vit}}\mathcal{L}_{\text{BCE}}$;
19
// Phase 2: Hierarchical RL Training. Train meta-controller and sub-controller through interaction with environment
20 for episode $e=1$ to $E$ do
21  Reset environment, observe initial state $s_{0}$;
22
 // Meta-controller loop: operates at coarse timescale (macro-steps)
23  for macro-step $t=1$ to $T_{\text{macro}}$ do
24     Capture visual input: $img_{t}\leftarrow\text{Camera}()$;
25     Predict blockage probabilities: $\hat{\mathbf{p}}_{t}\leftarrow\text{ViT}_{\theta_{vit}}(img_{t})$;
26     Form meta-state: $s_{t}^{h}=\{\hat{\mathbf{p}}_{t},\mathbf{L}_{t}\}$ where $\mathbf{L}_{t}=\{\mathbf{u}_{k,t}\}_{k=1}^{K}$ is UE locations;
27     Select subgoal (transmission mode): $g_{t}\leftarrow\pi_{h}(s_{t}^{h})+\mathcal{N}_{t}$ where $\mathcal{N}_{t}\sim\mathcal{OU}(0,\sigma_{o})$ is exploration noise;
28     Initialize cumulative reward for this macro-step: $R_{\text{meta}}\leftarrow 0$;
29
    // Sub-controller loop: operates at fine timescale (micro-steps)
30     for micro-step $\tau=1$ to $N_{\text{macro}}$ do
31        for $k=1$ to $K$ do
32           Extract geometric features from ray-tracing: $\mathbf{x}_{k}\leftarrow\{d_{k,n},\vartheta_{k,n},\psi_{k,n}\}_{n=1}^{N}$ via Eq. (15);
33           Estimate effective channel: $\hat{\mathbf{h}}_{\text{eff},k}\leftarrow\text{Transformer}_{\theta_{csi}}(\mathbf{x}_{k})$;
34
35        end for
36       Construct CSI matrix: $\hat{\mathbf{H}}_{\text{eff}}\leftarrow[\hat{\mathbf{h}}_{\text{eff},1},\ldots,\hat{\mathbf{h}}_{\text{eff},K}]$;
37
38       Form sub-controller state: $s_{\tau}^{l}=\{\hat{\mathbf{H}}_{\text{eff}},g_{t}\}$;
39        Select action with exploration noise: $a_{\tau}^{l}=\{\mathbf{W}_{\tau},\bm{\Theta}_{\tau}\}\leftarrow\pi_{l}(s_{\tau}^{l})+\mathcal{N}_{\tau}$ where $\mathcal{N}_{\tau}\sim\mathcal{OU}(0,\sigma_{o})$;
40
41       Apply beamforming matrix $\mathbf{W}_{\tau}$ at BS subject to power constraint $\mathbf{W}_{\tau}||_{F}^{2}\leq P_{\text{max}}$;
42        Apply RIS configuration $\bm{\Theta}_{\tau}$ at RIS;
43        Compute instantaneous sum SE: $r_{\tau}^{l}=\sum_{k=1}^{K}\text{SE}_{k,\tau}$ where $\text{SE}_{k,\tau}=\log_{2}(1+\text{SINR}_{k,\tau})$;
44        Observe next sub-controller state: $s_{\tau+1}^{l}$;
45
46       Store transition in replay buffer: $(s_{\tau}^{l},a_{\tau}^{l},r_{\tau}^{l},s_{\tau+1}^{l})\rightarrow\mathcal{B}_{l}$;
47        if $|\mathcal{B}_{l}|\geq B$ and $\tau\mod 4=0$ then
48           Sample random mini-batch of size $B$ from $\mathcal{B}_{l}$;
49           Compute target Q-value using target networks $Q_{l}^{\prime}$ and $\pi_{l}^{\prime}$;
50           Update sub-controller critic: $Q_{l}\leftarrow Q_{l}-\eta\nabla_{Q_{l}}\mathcal{L}_{Q}$ where $\mathcal{L}_{Q}$ is the TD error;
51
52          Update sub-controller actor: $\pi_{l}\leftarrow\pi_{l}+\eta\nabla_{\pi_{l}}Q_{l}(s,\pi_{l}(s))$;
53           Soft update target networks: $Q_{l}^{\prime}\leftarrow\tau Q_{l}+(1-\tau)Q_{l}^{\prime}$, $\pi_{l}^{\prime}\leftarrow\tau\pi_{l}+(1-\tau)\pi_{l}^{\prime}$;
54
55        end if
56
57       Accumulate with discount factor $\gamma_{l}$: $R_{\text{meta}}\leftarrow R_{\text{meta}}+\gamma_{l}^{\tau}r_{\tau}^{l}$;
58
59     end for
60
61    Observe next meta-state: $s_{t+1}^{h}$;
62     Store meta-transition in replay buffer: $(s_{t}^{h},g_{t},R_{\text{meta}},s_{t+1}^{h})\rightarrow\mathcal{B}_{h}$;
63     if $|\mathcal{B}_{h}|\geq B$ then
64        Sample random mini-batch of size $B$ from $\mathcal{B}_{h}$;
65        Compute target Q-value using target networks $Q_{h}^{\prime}$ and $\pi_{h}^{\prime}$ with discount factor $\gamma_{h}$;
66        Update meta-controller critic: $Q_{h}\leftarrow Q_{h}-\eta\nabla_{Q_{h}}\mathcal{L}_{Q}$;
67        Update meta-controller actor: $\pi_{h}\leftarrow\pi_{h}+\eta\nabla_{\pi_{h}}Q_{h}(s,\pi_{h}(s))$;
68        Soft update target networks: $Q_{h}^{\prime}\leftarrow\tau Q_{h}+(1-\tau)Q_{h}^{\prime}$, $\pi_{h}^{\prime}\leftarrow\tau\pi_{h}+(1-\tau)\pi_{h}^{\prime}$;
69
70     end if
71
72  end for
73
74 end for
Return: Trained models $\theta_{csi},\theta_{vit},\pi_{h},\pi_{l}$
Algorithm 1 Dual-Transformer HDRL Training Procedure
III-C Two-Timescale HDRL for Joint Optimization
The HDRL consists of two interacting agents: a Meta-controller operating at the coarse timescale, and a Sub-controller operating at the fine timescale. We detail the specific roles and optimization objectives of these components below:
Meta-controller (High-Level): This agent acts as the strategic planner. It utilizes the predictions of ViT to set a binary subgoal $g_{t}\in\{0,1\}^{K}$, effectively deciding the topology of the network (LoS vs. RIS-Relay) before the physical channel degrades.
Sub-controller (Low-Level): This agent acts as the real-time optimizer. It receives the instantaneous CSI estimates for all users, denoted as the matrix $\hat{\mathbf{H}}_{\text{eff},t}=[\hat{\mathbf{h}}_{\text{eff},1,t},\ldots,\hat{\mathbf{h}}_{\text{eff},K,t}]\in\mathbb{C}^{N\times K}$, where each column $\hat{\mathbf{h}}_{\text{eff},k,t}$ is the estimated effective channel for UE $k$ at time $t$ obtained from the CSI transformer. Its task is to execute the strategy by determining the optimal values for $\mathbf{W}$ and $\bm{\Theta}$ subject to the subgoal $g_{t}$ set by the meta-controller. This hierarchical decomposition builds upon recent advances in multi-timescale reinforcement learning for wireless networks [zhou2022hierarchical, habib2025llm], extending these concepts to the unique challenges of near-field RIS-assisted systems.
We model the decision-making process of these controllers as two interconnected Markov Decision Processes (MDPs), each defined by a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{R},\gamma\rangle$, as detailed below:
1) Meta-Controller MDP:
  • State Space ($\mathcal{S}_{h}$): The state $s_{t}^{h}\in\mathcal{S}_{h}$ includes the predicted blockage probabilities and user locations: $s_{t}^{h}=\{\hat{\mathbf{p}}_{t},\mathbf{L}_{t}\}$, where $\mathbf{L}_{t}=\{\mathbf{u}_{k,t}\}_{k=1}^{K}$ is the set of all UE positions at time $t$.
  • Action Space ($\mathcal{A}_{h}$): The action $g_{t}\in\mathcal{A}_{h}$ is a discrete binary vector $g_{t}\in\{0,1\}^{K}$, where $g_{t,k}=1$ instructs the sub-controller to use the RIS-assisted path for user $k$.
  • Reward Function ($\mathcal{R}_{h}$): The reward $R_{t}^{h}$ is the cumulative sum of the discounted rewards from the sub-controller over the macro-step: $R_{t}^{h}=\sum_{\tau=0}^{N_{\text{macro}}-1}\gamma_{l}^{\tau}r_{t+\tau}^{l}$, where $\tau\in\{0,1,\ldots,N_{\text{macro}}-1\}$ is the micro-step index within the current macro-step $t$, $r_{t+\tau}^{l}$ is the instantaneous reward at micro-step $\tau$, $N_{\text{macro}}$ is the number of micro-steps per macro-step, and $\gamma_{l}$ is the discount factor of the sub-controller.
2) Sub-Controller MDP:
  • State Space ($\mathcal{S}_{l}$): The state $s_{t}^{l}\in\mathcal{S}_{l}$ is composed of the fine timescale CSI matrix and the subgoal: $s_{t}^{l}=\{\hat{\mathbf{H}}_{\text{eff},t},g_{t}\}$.
  • Action Space ($\mathcal{A}_{l}$): The action $a_{t}^{l}\in\mathcal{A}_{l}$ consists of the continuous BS beamfocusing and RIS phase-shift matrices: $a_{t}^{l}=\{\mathbf{W}_{t},\bm{\Theta}_{t}\}$.
  • Reward Function ($\mathcal{R}_{l}$): The instantaneous reward $r_{t}^{l}$ is the sum SE as defined in Eq. (12): $r_{t}^{l}=\sum_{k=1}^{K}\text{SE}_{k,t}$, where $\text{SE}_{k,t}=\log_{2}(1+\text{SINR}_{k,t})$ is the achievable SE computed using the ground-truth channel state at time $t$.
We employ hierarchical DDPG [lillicrap2015continuous] to solve these MDPs. The meta-controller updates its policy based on cumulative rewards from execution of the sub-controller, learning to associate visual blockage threats with long-term strategic mode switches. The critic networks are trained using temporal difference (TD) learning, with separate experience replay buffers for each hierarchy level to ensure stable convergence.
III-D Complexity Analysis
We analyze the computational complexity of our proposed approach relative to standard methods. Traditional channel estimation for RIS-assisted systems requires estimating the full cascaded channel (BS-RIS-UE). Standard methods such as Least Squares (LS) or MMSE scale with the product of the array sizes, yielding complexity $\mathcal{O}(KMN)$ or even $\mathcal{O}(KM^{2})$ depending on the pilot scheme [hu2021two], where $M$ is the number of RIS elements. This dependence on $M$ becomes a bottleneck for large intelligent surfaces.
In contrast, our proposed framework decouples the dimensionality. The CSI transformer operates on ray-tracing features at the BS, achieving complexity $\mathcal{O}(L_{\text{csi}}(N^{2}+Nd)K)$ per inference. Notably, this is independent of $M$, as the model learns to infer the effective link quality directly from BS-side geometry. The dependence on $M$ is handled solely by the action head of HDRL agent. The inference complexity of the agent is $\mathcal{O}(N_{\text{layer}}d_{\text{hidden}}^{2}+NKd_{\text{hidden}}+(N+M)d_{\text{hidden}})$, where the final term $(N+M)$ accounts for the joint optimization of BS beamforming and RIS phase shifts.
Furthermore, the hierarchical design amortizes the heavy ViT computation over the macro-timescale $T_{\text{macro}}$. The complete training procedure for the DT-HDRL framework is formalized in Algorithm 1.
IV Numerical Results and Discussion
IV-A DeepVerse 6G Dataset Description
We employ the O1 Urban scenario from the DeepVerse 6G ray-tracing dataset [demirhan2025deepverse], a publicly available resource specifically designed for machine learning-driven 6G research. The O1 scenario captures a dense metropolitan environment, featuring high-rise buildings with complex façade materials (glass, concrete, metal) that realistically model reflection, scattering, and diffraction at mmWave/THz frequencies. The dataset provides synchronized wireless channel data and visual imagery, enabling joint training of our dual-transformer framework. It contains channel realizations generated using a GPU-accelerated ray-tracing engine with up to 5th-order reflections and diffuse scattering. Multiple RGB cameras are mounted on the BS, covering a wide field-of-view to capture the environmental dynamics. The inter-site distances place UEs firmly in the NFC region for large antenna arrays, as validated by the Rayleigh distance calculation in Table I.
To generate ground-truth blockage labels for ViT training, we leverage the synchronized mmWave channel measurements and ray-tracing data. For each time instance and UE, we extract the LoS path status directly from the ray-tracing simulator, which tracks line-of-sight (LoS) visibility between the BS and each UE by computing ray intersections with environmental objects. The binary blockage label $b_{k}\in\{0,1\}$ for UE $k$ at each time step is derived as:
$b_{k}=\begin{cases}1,&\text{if }PL_{\text{LoS},k}<PL_{\text{th}}\\ 0,&\text{otherwise}\end{cases}$ (26)
where $PL_{\text{LoS},k}$ is the measured LoS path loss from the mmWave channel data, and $PL_{\text{th}}$ is the threshold specified in Table I. The labeling process iterates through all time indices, BS indices, and UE indices in the dataset. This automated procedure generates synchronized data for supervised ViT training, ensuring that visual features are correctly associated with wireless channel conditions.
IV-B Simulation Setup
We evaluate the performance of our proposed DT-HDRL using the O1 urban scenario from the DeepVerse 6G ray-tracing dataset [demirhan2025deepverse]. The system comprises a downlink multi-user MISO configuration with $K=10$ single-antenna UEs distributed within a $50\times 50\times 1$ m3 region. The BS employs an XL-MIMO UPA with $N=1024$ antennas configured as $N_{y}=32\times N_{z}=32$ at $0,0,0]^{T}$, while the RIS with $M=100$ elements is deployed at $15,0,15]^{T}$ m. Therefore, the BS-RIS distance is about 21.2 m, which is clearly inside the Rayleigh distance, which is 82 m. Camera arrays and sensing devices are co-located at the BS for synchronized visual and channel data acquisition. Key parameters are summarized in Table I.
TABLE I: Simulation Parameters
Parameter Value
System Configuration
BS Antennas ($N=N_{y}\times N_{z}$) $1024=32\times 32$
RIS Elements ($M=M_{y}\times M_{z}$) $100=10\times 10$
Carrier Frequency ($f_{c}$) 3.5 GHz
Wavelength ($\lambda$) 85.7 mm
Effective Aperture ($A_{\text{eff}}$) $\lambda^{2}/4\pi$
System Bandwidth 100 MHz
Noise Power Spectral Density ($\sigma^{2}$) -94 dBm
SNR Range -10 to 50 dB
Number of UEs ($K$) 10
UE Antennas 1
BS Element Spacing ($d_{\text{BS}}$) $\lambda/2$
RIS Element Spacing ($d_{\text{RIS}}$) $\lambda/5$
BS Position ($\mathbf{u}_{\text{BS}}$) $0,0,0]^{T}$ m
RIS Position ($\mathbf{u}_{\text{RIS}}$) $15,0,15]^{T}$ m
UE Region $50\times 50\times 1$ m3
Rayleigh Distance ($Z_{R}$) 82 m
Maximum Transmit Power ($P_{\text{max}}$) 35 dBm
Minimum SE per User ($\text{SE}_{\text{min}}$) 1 bps/Hz
Visual Data Collection
Camera Frame Rate 6.5 samples/s
Image Resolution $960\times 540$ pixels
Frames per Sequence ($F$) 10
Blockage Threshold ($PL_{\text{th}}$) Noise floor + 10 dB
Timescale Configuration
Meta-Controller Timescale 154 ms
Sub-Controller Timescale ($T$) 1 ms
Micro-steps per Macro ($N_{\text{macro}}$) 154
Training Configuration
Number of Training Episodes ($E$) 5000
Macro-steps per Episode ($T_{\text{macro}}$) 100
Training Samples ($N_{\text{data}}$) 2,000
Training Epochs 200 (CSI), 1000 (ViT/HDRL)
Learning Rate ($\eta$) $10^{-4}$
Batch Size ($B$) 64 (CSI), 256 (HDRL)
Optimizer Adam
BCE Numerical Stability Constant ($\epsilon$) $10^{-7}$
HDRL Parameters
Replay Buffer Size Sub ($|\mathcal{B}_{l}|$) $10^{5}$
Replay Buffer Size Meta ($|\mathcal{B}_{h}|$) $10^{4}$
Discount Factor Sub ($\gamma_{l}$) 0.99
Discount Factor Meta ($\gamma_{h}$) 0.9
Target Network Update Rate ($\tau$) 0.001
Exploration Noise Std ($\sigma_{o}$) 0.3 (Ornstein-Uhlenbeck)
DDPG Update Frequency Every 4 steps
CSI Transformer Architecture
Embedding Dimension ($d_{\text{model}}$) 512
Number of Layers ($L$) 6
Number of Attention Heads ($h$) 8
Key/Query Dimension ($d_{k}$) 64
Value Dimension ($d_{v}$) 64
Activation Function ReLU
ViT Architecture
Embedding Dimension ($D$) 768
Number of Encoder Blocks ($L_{\text{ViT}}$) 12
Number of Attention Heads 12
Patch Size ($P$) $16\times 16$ pixels
Number of Patches ($N_{p}$) 2025
MLP Hidden Layers [512, 256, 128]
Activation Function ReLU
IV-C Baseline Methods
To evaluate the performance of our proposed DT-HDRL framework, we compare against the following baseline methods:
1) CNN-based CSI Estimation: A convolutional neural network with three convolutional layers (64, 128, 256 filters) followed by fully connected layers for channel estimation. This baseline represents traditional deep learning approaches that rely on local receptive fields.
2) Transformer for Blockage Prediction: A vanilla transformer encoder without patch-based processing, serving as a comparison to our ViT architecture for blockage prediction tasks.
3) DRL without RIS (w/o RIS, DRL): A single-level deep reinforcement learning agent (DDPG) that directly optimizes BS beamforming without hierarchical decomposition and without RIS assistance.
4) DRL with RIS (w/ RIS, DRL): A single-level DDPG agent that jointly optimizes BS beamforming and RIS phase shifts without timescale decomposition.
5) HDRL without RIS (w/o RIS, HDRL): A hierarchical DDPG agent with dual-timescale decomposition (meta-controller for mode selection and sub-controller for beamforming optimization) but without RIS assistance, demonstrating the isolated contribution of the hierarchical structure.
All baseline methods use the same system parameters and are trained with identical hyperparameters for fair comparison.
IV-D Convergence and CSI Estimation Accuracy
We first validate the effectiveness of our transformer-based CSI estimator by comparing it against a CNN baseline, which represents the current state-of-the-art in learning-based channel estimation. This comparison isolates the contribution of the self-attention mechanism in capturing near-field spatial dependencies. Subsequently, we evaluate the complete DT-HDRL system performance, including both transformers and hierarchical control.
Training Convergence: Fig. 2 compares the training loss (MSE). The transformer achieves a 73% reduction in final MSE compared to the CNN. This gain is attributed to the ability of the self-attention mechanism to model the non-local correlations of spherical waves across the large aperture, whereas CNNs are limited by the local receptive fields of their kernels.
SNR Performance: As shown in Fig. 3, the CSI transformer maintains strong performance across the entire SNR range. While the CNN degrades sharply at low SNRs, the attention mechanism of transformer effectively filters noise by focusing on dominant propagation paths, ensuring robust input for the beamfocusing optimization.
Figure 2: Training and validation loss comparison. The transformer achieves 73% lower final MSE with stable convergence.
Figure 3: Normalized MSE for different received SNR. The CSI transformer maintains low estimation error even in noisy conditions.
IV-E Blockage Prediction Performance
Fig. 4 compares the training loss convergence over 1000 epochs for ViT and another Transformer architecture for blockage prediction, showing that ViT achieves a lower and more stable final loss, while both models converge rapidly in the initial epochs. The superior performance of ViT stems from its patch-based processing and positional encodings, which are well-suited for capturing spatial motion patterns in video sequences.
The ViT blockage predictor achieves an F1-score of 0.92 (Table II). The model correctly predicts blockages 5 frames before occurrence ($\approx 769$ ms). This future window is the key enabler for the HDRL meta-controller, allowing it to reconfigure the RIS proactively rather than reacting to a link failure.
Figure 4: Training loss convergence for blockage prediction. ViT achieves lower final loss with slightly faster initial convergence.
TABLE II: Blockage Prediction Performance
Method Precision Recall F1-Score
Proposed Method 0.94 0.90 0.92
Transformer Baseline 0.86 0.82 0.84
IV-F SE Performance
Dynamic Evolution: Fig. 5 tracks the average SE over the training steps. Our proposed DT-HDRL framework achieves a converged average SE of 285.92 bps/Hz, significantly outperforming the DRL baseline without RIS, which saturates at 242.65 bps/Hz. This represents a performance gain of approximately 18%. The superiority over the baseline highlights the value of timescale decomposition: the meta-controller stabilizes the learning process by abstracting the blockage dynamics, allowing the sub-controller to focus purely on fast-fading compensation.
Impact of Transmit Power: Fig. 6 shows that at $P_{\text{max}}=35$ dBm, our framework outperforms the non-RIS, DRL baseline by $\approx 9.5$ bps/Hz. This confirms that the HDRL agent effectively utilizes the blockage prediction results and RIS as a passive beamformer to extend coverage to edge users, converting the additional transmit power into useful signal gain rather than interference.
Figure 5: Average accumulative reward (sum SE) for different simulation steps.
Figure 6: Accumulative SE for different BS transmit power.
Ablation Study (Theoretical): It is worth noting the contribution of the specific geometric features used in our CSI transformer. Standard far-field estimators rely solely on Angle of Arrival (AoA). In the NFC, AoA is distance-dependent. By explicitly embedding the distance $d_{k,n}$ as part of the input feature vector $\mathbf{x}_{k}$ (Eq. 15), our model linearizes the phase wavefront in the high-dimensional latent space. Without this distance embedding, the attention mechanism would fail to distinguish between users at the same angle but different depths, leading to significant “beam squint” (frequency-dependent beam deviation across subcarriers) losses [gao2023wideband]. The proposed architecture effectively acts as a non-linear phase compensator, validating its superiority over standard CNNs shown in Fig. 2.
IV-G Array Scaling Analysis
XL-MIMO Scaling: Fig. 7 shows performance for different BS antenna count $N$. At $N=256$, our HDRL with RIS achieves approximately 57 bps/Hz, outperforming the DRL baseline (52 bps/Hz) by roughly 5 bps/Hz. The performance gap between our method and the baselines widens with larger arrays, indicating that the transformer-based CSI estimator scales more effectively to high-dimensional spaces than standard estimators, preserving the beamfocusing gain of ELAA systems.
RIS Element Scaling: Fig. 8 illustrates the impact of RIS size $M$. Increasing $M$ from 25 to 900 yields significant gains, with our method reaching $\approx 59$ bps/Hz compared to $\approx 54$ bps/Hz for the baseline without RIS. However, we observe a performance saturation beyond $M=625$. This suggests a physical limit imposed by the spatial correlation of the channel; simply adding more elements beyond the coherence distance yields diminishing returns, a crucial insight for practical RIS dimensioning.
Figure 7: Average SE for different numbers of BS antennas NN, showing consistent gains from the proposed HDRL with RIS across all array sizes.
Figure 8: Average SE for different numbers of RIS reflecting elements, demonstrating scaling benefits with increased RIS size.
IV-H Computational Complexity Validation
We validate the practical feasibility of our framework by measuring inference time and computational complexity on a workstation equipped with an Intel Core i7-11370H CPU, 24 GB RAM, and an RTX3060 GPU with specific system parameters: $N=128$, $M=100$, and $K=10$. This hardware setup represents a realistic deployment scenario for edge computing in 6G networks, and similar complexity validation approaches have been adopted in recent works on transformer-based wireless systems [jin2025near, yellapragada2025computationally]. As shown in Table III, the total computational cost per frame is 5.09 GFLOPs. This is dominated by the ViT blockage predictor (4.82 GFLOPs) operating at the coarse timescale (once per 154 ms). The CSI transformer ($L=6,d=512$) and HDRL agent ($L=3,d_{\text{hidden}}=512$) together require only 0.27 GFLOPs per TTI. Given that an RTX 3060 offers over 12 TFLOPs of compute, the required throughput is well within real-time limits at the camera sampling rate of 6.5 samples/s.
TABLE III: Computational Complexity per Frame
Component GFLOPs
CSI Transformer ($N=128$, $K=10$) 0.12
ViT Blockage Predictor ($p=2025$, $L=12$) 4.82
HDRL Agent (Optimizing $N=128$, $M=100$) 0.15
Total per Frame 5.09
V Conclusion
The proposed DT-HDRL framework demonstrates substantial performance improvements in RIS-assisted NFC system. The CSI transformer achieves significant reduction in estimation error compared to CNN baselines, while the ViT blockage predictor provides sufficient advance warning to enable proactive reconfiguration. The hierarchical control structure effectively manages dual timescales, with the meta-controller making strategic routing decisions at coarse intervals while the sub-controller performs continuous beamforming and phase shift optimization. Extensive evaluations across varying transmit power, antenna array sizes, and RIS configurations demonstrate consistent SE improvements. The framework maintains real-time feasibility on a GPU hardware. By aligning learning architectures with the inherent physics of wireless propagation and environmental dynamics, this work opens a pathway toward intelligent, proactive 6G networks capable of adaptive near-field beamfocusing and blockage mitigation.
Acknowledgment
This work has been supported by the NSERC Canada Research Chairs program, MITACS, and Ericsson. We also wish to commemorate our co-author, Han, whose exceptional commitment and meaningful contributions strengthened this work. Her warmth and enthusiasm will be sincerely missed.